A baker made 20 pies. A boy scout troop buys 1/4 of his pies, apreschool teacher buys 1/3 of his pies, and a caterer buys 1/6 of his pies. How many pies does the baker have left?
Answer by oberobic(2304)

1/4 + 1/3 + 1/6 = ?
The common denominator is 12.
1/4 = 3/12
1/3 = 4/12
1/6 = 2/12
.
9/12 have been sold = 75%.
.
So 25% of the pies have not been sold.
25% of 20 = 1/4 * 20 = 5
.
The baker still has 5 pies.
.
But there is a problem with this problem statement:
You cannot buy 1/3 of 20 pies unless the baker lets you buy 6 2/3 pies.
Similarly, you cannot buy 1/6 of 20 pies unless the baker lets you buy 3 1/3 pies.
The math works, but I don't know any bakers who sell partial pies.
.
Done.